# Simpler Diffusion (SiD2): 1.5 FID on ImageNet512 with pixel-space diffusion

### v1

@@ 2.4 Distillation
Sampling from a pixel-space diffusion model naively often requires more function evaluations than sampling from a latent-space diffusion model. For example, the popular Stable Diffusion model (Rombach et al., 2022) is usually sampled using 20-50 sampling steps, while pixel-space models are sampled with 100+ steps (Hoogeboom et al., 2023). It would therefore seem that latent models are more efficient at test time. However, recent advances in the distillation of diffusion models show that pixel-space models can be easily adapted into few-step samplers that maintain high sample quality (Salimans & Ho, 2022; Xiao et al., 2021; Kim et al., 2024a; Yin et al., 2023; Xu et al., 2023; Luo et al., 2024). In this work, we use the moment matching distillation method proposed by Salimans et al. (2024) to distill our pixel-space models for use with 8-16 sampling steps, which empirically we find maintains or even improves the sample quality of our models (see Section 4).
@@ 3 Related Work
In literature the first end-to-end diffusion work that directly generated high resolution images was (Dhariwal & Nichol, 2021). Given the poor performance of end-to-end models, two popular alternatives were introduced: cascaded diffusion (Ho et al., 2022) and latent diffusion (Rombach et al., 2022). Significant improvements were made to end-to-end diffusion by shifting noise schedules (Hoogeboom et al., 2023; Chen, 2023) and subsequently loss weightings (Kingma & Gao, 2023). Other approaches such as multi-resolution losses (and possibly architecture adaptions) (Gu et al., 2022; 2023; Hoogeboom et al., 2023) have also been introduced with reasonable successes. However, the current best known latent model EDM2 (Karras et al., 2023) outperforms these prior end-to-end approaches by a large margin. In contrast, our models are competitive with latent models, outperforming DiT-XL/2 (Peebles & Xie, 2022) and only performing slightly worse than EDM2.
@@
The U-Net architecture is a fully convolutional network with long-range skip-connections introduced for (medical) image segmentation (Ronneberger et al., 2015; Shelhamer et al., 2016). The main idea of U-Net is to combine a contracting path (which encodes the image semantics) with long skip-connections at the same resolution to enable precise localization.
With the introduction of ViTs (Dosovitskiy et al., 2021) as an alternative for convolutional layers, the U-Net has been converted to U-ViTs using a similar network but replacing (some) ResNet blocks with ViTs, either with spatial down-sampling (Hoogeboom et al., 2023) or without (Bao et al., 2023). Long-range skip-connections play an pivotal role in these networks, for example for fine-tuning controllable diffusion models (Jiang et al., 2023) and for the stability and training speed (Huang et al., 2023; Ma et al., 2024). In these latter works, it is shown that multiplying the skip-connections with a (constant) scalar ($<1$) improves training stability and speed, because of the propagation of noise throughout the network. In our work, we reduce the number of skip-connections significantly and initialise the network such that it resembles an identify function.
@@
Fine-tuning on a higher resolution at large scale was effectively done in Stable Diffusion (Rombach et al., 2022) to limit the training cost, which is a form of flop heavy scaling over different resolutions. In this paper we show that flop heavy scaling is also preferred from a regularization perspective, which can be done either by finetuning from a lower resolution, or by lowering the patching size.
4 Experiments
4.1 Sigmoidal weighting functions versus others
Loss weighting
In this experiment, we want to understand the effect of loss weighting on the final image quality of a model. For the noise schedule, we use an interpolated cosine schedule because it has relatively broad coverage over logsnr values. We compare the weightings corresponding to the original ${\bm{{\epsilon}}}$-mse (Ho et al., 2020) and the multiscale ${\bm{{\epsilon}}}$-mse (Hoogeboom et al., 2023), the EDM shifted weightings (Karras et al., 2022), including the original and monotonic versions (Karras et al., 2022; Kingma et al., 2021), and the proposed sigmoidal weighting with shift $b=-3$. As can be seen in Figure 5, the sigmoidal weighting outperforms all other weightings, with EDM monotonic shifted being the closest alternative. Appendix B contains additional experiments, where we also evaluate the sigmoidal weighting against the closest competitor without guidance. Here again sigmoidal weighting is consistently better. We also evaluate a time-shifted variant, where the bias depends on the progress in training.
Training efficiency of noise schedules
Since the training noise schedule can be decoupled from the loss weighting and it only affects the training efficiency but not the final performance, one can measure the convergence speed over the training process caused by leveraging particular noise schedules. In this experiment we test the cosine schedule (Ho et al., 2020), the shifted cosine and the interpolated cosine schedules (Hoogeboom et al., 2023). For fair comparison, we fix the sampling noise schedule as the cosine interpolated schedule. Interestingly, when the sigmoidal weighting is used, the choice of noise schedule does not affect the training efficiency significantly. Arguably, the shifted schedule is a bit more sample efficient initially but all schedules perform similarly towards the end. This suggests that there is a lot of transfer across different logsnr values. We also tried the EDM training noise schedule with the sigma power sampling noise schedule (Karras et al., 2022) (not plotted), with both shifted or the original version, but its narrower logsnr sampling schedule led to performance degradation, achieving about 9 FID at best. This suggests that for high resolution, a wide logsnr range is crucial to cover the important noise levels.
Bias over resolutions
In Table 1, we show the effect of the bias term $b$ on the final performance when trained using a flop heavy model. All models use the same guidance and similar guidance intervals, where the lower bound of the guidance interval has been shifted according to resolution (again approximately 1.5 per 2x resolution change), see Appendix A for more details..
Sigmoidal weighting versus a multiscale variant
In literature, high resolution generation can greatly benefit from having a multiscale loss (Gu et al., 2022; 2023; Hoogeboom et al., 2023). Inspired by that, we examine the sigmoidal weighting loss against a multiscale variant of it. To that end we create a multiscale variant that we name the power loss. In addition to the sigmoidal weighting that effectively upweights low frequency components of the data and downweights high frequency components, the power loss aims to amplify the loss of low sub-band (smooth) signals in the spatial domain via Haar wavelets. For details on this loss see Appendix A.1.
In Table 2 we see that the bias should be adjusted based on the number of Haar transform downsampling levels. Despite being competitive, the performance of power loss is roughly on par with the one with sigmoid weighting but without multiscaling (i.e. downsample levels $=0$), if we tune the bias of the sigmoidal weighting more aggressively. Interestingly, lowering the bias has almost the same effect as increasing the scale. Therefore, building the multiscale variant is not necessary with the sigmoidal weighting.
Figure 5: Effect of weighting functions on the FID score of ImageNet512 over the training iteration.
Figure 6: Effect of training noise schedules on the FID score of ImageNet512 over the training iteration. Fortunately with the sigmoidal weighting with $b$ $=$ $-3$, noise schedules have little effect on the training process.
Table 1: Sigmoidal weighting performance over different resolutions and shifts with the flop heavy architecture. Larger resolutions require more negative bias.
@@ 4.1 Sigmoidal weighting functions versus others
Table 2: Power loss performance over downsample levels and bias on ImageNet512 with the small model. Increasing downsampling has a similar effect as decreasing bias.
@@ 4.2 Scaling channels vs tokens on ImageNet
In this section we show the difference between scaling a model 4x FLOPS, either via channels or by creating larger resolution feature maps (or put differently, feature maps with a higher token count). Note that our small model on ImageNet512 is patched 4 x 4. This model is scaled either by increasing channels by two (named model size 2x), or by reducing patching to 2 x 2 (named flop heavy, because FLOPS over params is much higher). Indeed as can be seen in Figure 8, the flop heavy model performs much better than the model size 2x, when trained in a standard way on ImageNet512. Note that unlike most experiments, in these experiments it was necessary to use very different optimal guidance settings, more details in Appendix A.
We confirm this is an issue related to overfitting, by training the model size 2x with distribution augmentation (Jun et al., 2020) the model size 2x variant performs much better, with an FID of 1.6 instead of 1.9 (latter in Figure 8). Still, without any augmentation beyond the standard horizontal flips, the flop heavy model achieves a score of 1.54.
Table 3: Literature Comparison on ImageNet.
Method FID
@@
RIN (Jabri et al., 2022) 2.75
SD, U-ViT-L (Hoogeboom et al., 2023) 1.94
StyleGAN-XL (Sauer et al., 2022) 1.81
VDM$+$ (Kingma & Gao, 2023) 1.75
DisCo-Diff (Xu et al., 2024) 1.73
PaGoDA (Kim et al., 2024b) 1.48
SiD2, Flop Heavy (ours) 1.26
@@
ADM-G (Dhariwal & Nichol, 2021) 4.59
ADM-U (Dhariwal & Nichol, 2021) 3.94
CDM (Ho et al., 2022) 4.88
RIN (Jabri et al., 2022; Chen, 2023) 3.52
SD, U-ViT-L (Hoogeboom et al., 2023) 2.44
MDM (Gu et al., 2023) 3.51
StyleGAN-XL (Sauer et al., 2022) 2.30
DiT-XL/2 (Peebles & Xie, 2022) 2.27
VDM++ (Kingma & Gao, 2023) 2.12
PaGoDA (Kim et al., 2024b) 1.56
SiD2, small (ours) 1.72
SiD2, Flop Heavy (ours) 1.38
@@
ADM-G (Dhariwal & Nichol, 2021) 7.72
RIN (Jabri et al., 2022; Chen, 2023) 3.95
ADM-U (Dhariwal & Nichol, 2021) 3.85
LDM-4-G (Rombach et al., 2022) 3.60
SD, U-ViT-L (Hoogeboom et al., 2023) 3.02
DiffiT (Hatamizadeh et al., 2023) 2.67
StyleGAN-XL (Sauer et al., 2022) 2.40
DiT-XL/2 (Peebles & Xie, 2022) 3.04
VDM++ (Kingma & Gao, 2023) 2.65
MAGVIT-v2 (Yu et al., 2024) 1.91
PaGoDA (Kim et al., 2024b) 1.80
EDM2-S (interval) (Kynkäänniemi et al., 2024) 1.68
EDM2-XXL (interval) (Kynkäänniemi et al., 2024) 1.40
SiD2, small (ours) 2.19
SiD2, Flop Heavy (ours) 1.48
SiD2, Flop Heavy 16-step distilled (ours) 1.50
@@ 4.3 Removing block skip connection
In this section we study the difference between the residual skip-connections and the blockwise skip-connections. As hypothesized, the small model performance drops a bit because of this. For ImageNet512 a deficit of about 0.2 FID (in Figure  9) and for Imagenet256 a deficit of about 0.05 FID. That the task of fitting ImageNet256 is less affected by removing skip-connections can be explained by the task being easier: The task of fitting ImageNet256 is a subset of fitting ImageNet512.
@@ 4.4 Literature comparison
Here we compare different models from the literature. In the first section end-to-end diffusion models are compared, and in the second section we compare between all types of generative models. In Table 3 it can be seen that our SiD2 is very competetive with existing generative models: On ImageNet128 and ImageNet256 it achieves state-of-the-art performance, on ImageNet512 it is a little worse but still competitive with the best latent diffusion model and outperforms all other models, end-to-end and latent.
@@
The training cost SiD2 is another helpful way to see how the model scales, as depicted in Figure 7). The most important thing is that the performance of our end-to-end approach outperforms other significantly. Observe for example the performance of our small variant (using no guidance interval to keep this comparison fair), which roughly the same performance as the VDM++ while only using a third of the training cost, and a fourth of the FLOPS in a single forward pass (see Figure 11 in Appendix A). It is also competitive with the best currently known latent diffusion approach EDM2 (Karras et al., 2023), but in fairness EDM2 scales a little bit better. These models attain even better performance for a smaller training cost, and a similar story holds for forward pass complexity. End-to-end diffusion can work much better than previously demonstrated, but it is still somewhat outperformed by latent diffusion on quality versus compute comparisons.
0.00.00.00.00.50.50.50.51.01.01.01.01.51.51.51.52.02.02.02.02.52.52.52.53.03.03.03.03.53.53.53.5111122223333555510101010ADMADM-UDiT-XL/2 (interval)RINSD, U-ViT-LVDM++EDM2-S (interval)EDM2-XXL (interval)EDM2-SEDM2-LEDM2-XXL(no interval)Small(no interval)Flop HeavyFlop Heavy (finetune)Training cost (zettaflops)Previous (latent)PreviousSimpler diffusion (ours)
Figure 7: FID versus training cost on ImageNet512 between other end-to-end diffusion models, and several latent diffusion models. Note that the EDM-2 training curves are without guidance, but the best performance with guidance intervals is reported with a separate marker. The figure source adapted has been adapted from EDM2 (Karras et al., 2023). Following their assumptions, one training iteration is three times as expensive as evaluating the model. Our Flop Heavy (finetune) is trained by first training a small model on ImageNet256 for 400k iterations and then finetuning the same model on ImageNet512.
@@
Table 4: Zero-shot text to image performance on the MSCOCO dataset.
[…]
eDiff-I (Balaji et al., 2022) 256 6.95
Parti (Yu et al., 2022) 256 7.23
Imagen (Saharia et al., 2022) 256 7.27
simple diffusion (Hoogeboom et al., 2023) 256 8.3
simple diffusion (Hoogeboom et al., 2023) 512 9.6
SDv1.5 (eval DMD2) (Rombach et al., 2022) 512 7.21
DMD2 (Yin et al., 2024) 512 8.35
SiD2 512 8.1
SiD2 16-step distilled 512 6.68
. Refer to caption

### v2

@@ 2.4 Distillation
Sampling from a pixel-space diffusion model naively often requires more function evaluations than sampling from a latent-space diffusion model. For example, the popular Stable Diffusion model [32] is usually sampled using 20-50 sampling steps, while pixel-space models are sampled with 100+ steps [14]. To reduce the number of sampling steps, we could use either more advanced solvers or distillation. Recently distillation methods show that pixel-space models can be adapted into few-step samplers that maintain high sample quality [36, 44, 23, 47, 45, 29]. In this work, we use moment matching distillation [37] which is currently one of the best distillation methods available in the few step regime. We distill our pixel-space models for use with 8-16 sampling steps, which empirically we find maintains or even improves the sample quality of our models.
Figure 5: Effect of weighting functions in the loss on the FID score of ImageNet512 over the training iterations. The proposed sigmoid weighting with $b\!=\!-3$ shows best performance consistent across all training iterations.
Figure 6: Effect of training noise schedules on the FID score of ImageNet512 over the training iterations. Fortunately, with our proposed sigmoid weighting ($b\!=\!-3$), the exact choice of the noise schedules has little effect on the training process.
@@ 3 Related Work
The first end-to-end diffusion work that directly generated high resolution images was [4]. Given the poor performance of end-to-end models, two popular alternatives were introduced: cascaded diffusion [12] and latent diffusion [32].
Cascaded diffusion approaches have been shown to scale to large datasets and models to generate high quality and high resolution imagery, albeit in multiple stages [12, 42, 52, 40], or with fine-tuning [43]. Although the focus of this paper has been single stage generation up to a resolution of $1024^{2}$, the loss weightings may also be used for cascaded models.
Significant improvements were made to end-to-end diffusion by shifting noise schedules [14, 3] and subsequently loss weightings [25]. Other approaches such as multi-resolution losses (and possibly architecture adaptions) [6, 7, 14] added marginal gains. However, the current best known latent model EDM2 [20] outperforms these prior end-to-end approaches by a large margin. In contrast, our models are competitive with latent models, outperforming DiT-XL/2 [31] and only performing slightly worse than EDM2.
@@
The U-Net architecture is a fully convolutional network with long-range skip-connections introduced for (medical) image segmentation [33, 39]. The main idea of U-Net is to combine a contracting path (which encodes the image semantics) with long skip-connections at the same resolution to enable precise localization.
With the introduction of ViTs [5] as an alternative for convolutional layers, the U-Net has been converted to U-ViTs using a similar network but replacing (some) ResNet blocks with ViTs, either with spatial down-sampling [14] or without [2]. Long-range skip-connections play an pivotal role in these networks, for example for fine-tuning controllable diffusion models [17] and for the stability and training speed [15, 30]. In these latter works, it is shown that multiplying the skip-connections with a (constant) scalar ($<1$) improves training stability and speed, because of the propagation of noise throughout the network. In our work, we reduce the number of skip-connections significantly and initialise the network such that it resembles an identify function.
@@
Fine-tuning on a higher resolution at large scale was effectively done in Stable Diffusion [32] to limit the training cost, which is a form of flop heavy scaling over different resolutions. In this paper we show that flop heavy scaling is also preferred from a regularization perspective, which can be done either by finetuning from a lower resolution, or by lowering the patching size.
Table 1: Sigmoid weighting performance (left) with the flop heavy architecture. Larger resolutions require more negative bias. Power loss performance (right) on ImageNet512 with the small model. Increasing downsampling has a similar effect as decreasing bias, hence we prefer the more simple sigmoid weighting.
(a) Sigmoid weighting
@@
(b) Power loss weighting
@@
0.00.00.00.00.50.50.50.51.01.01.01.01.51.51.51.52.02.02.02.02.52.52.52.53.03.03.03.03.53.53.53.5111122223333555510101010ADMADM-UDiT-XL/2 (interval)RINSD, U-ViT-LVDM++EDM2-S (interval)EDM2-XXL (interval)EDM2-SEDM2-LEDM2-XXL(no interval)Small(no interval)Flop HeavyFlop Heavy (finetune)Training cost (zettaflops)Previous (latent)Previous (pixel)Simpler diffusion (ours)
Figure 7: FID versus training cost on ImageNet512. Note that the EDM-2 training curves are without guidance, but the best performance with guidance intervals is reported with a separate marker, the figure source adapted has been adapted from EDM2 [20]. Following their assumptions, one training iteration is three times as expensive as evaluating the model. Our Flop Heavy (finetune) is trained by first training a small model on ImageNet256 for 400k iterations and then finetuning the same model on ImageNet512.
4 Experiments
4.1 Sigmoid weighting versus others
Loss weighting
In this experiment, we want to understand the effect of loss weighting on the final image quality of a model. For the noise schedule, we use an interpolated cosine schedule because it has relatively broad coverage over logsnr values. We compare the weightings corresponding to the original ${\bm{{\epsilon}}}$-mse [11] and the multiscale ${\bm{{\epsilon}}}$-mse [14], the EDM shifted weightings [19], including the original and monotonic versions [19, 26], and the proposed sigmoid weighting with shift $b=-3$. As can be seen in Figure 6, the sigmoid weighting outperforms all other weightings, with EDM monotonic shifted being the closest alternative. Appendix C contains additional experiments, where we also evaluate the sigmoid weighting against the closest competitor without guidance. Here again sigmoid weighting is consistently better. We also evaluate a time-shifted variant, where the bias depends on the progress in training.
Training efficiency of noise schedules
Since the training noise schedule can be decoupled from the loss weighting and it only affects the training efficiency but not the final performance, one can measure the convergence speed over the training process caused by leveraging particular noise schedules. In this experiment we test the cosine schedule [11], the shifted cosine and the interpolated cosine schedules [14] (Figure 6). For fair comparison, we fix the sampling noise schedule as the cosine interpolated schedule. Interestingly, when the sigmoid weighting is used, the choice of noise schedule does not affect the training efficiency significantly. Arguably, the shifted schedule is a bit more sample efficient initially but all schedules perform similarly towards the end. This suggests that there is a lot of transfer across different logsnr values. We also tried the EDM training noise schedule with the sigma power sampling noise schedule [19] (not plotted), with both shifted or the original version, but its narrower logsnr sampling schedule led to performance degradation, achieving about 9 FID at best. This suggests that for high resolution, a wide logsnr range is crucial to cover the important noise levels.
Bias over resolutions
In Table 1(a), we show the effect of the bias term $b$ on the final performance when trained using a flop heavy model. All models use the same guidance and similar guidance intervals, where the lower bound of the guidance interval has been shifted according to resolution (again approximately 1.5 per 2x resolution change), see Appendix B for more details.
Sigmoid weighting versus the power loss
In literature, high resolution generation can greatly benefit from having a multiscale loss [6, 7, 14]. To that end we create a multiscale variant that we name the power loss. In addition to the sigmoid weighting that effectively upweights low frequency components of the data and downweights high frequency components, the power loss aims to amplify the loss of low sub-band (smooth) signals in the spatial domain via Haar wavelets. For details on this loss see Appendix B.2.
In Table 1(b) we see that the bias should be adjusted based on the number of Haar transform downsampling levels. Despite being competitive, the performance of power loss is roughly on par with the one with sigmoid weighting but without multiscaling (i.e. downsample levels $=0$), if we tune the bias of the sigmoid weighting more aggressively. Interestingly, lowering the bias has almost the same effect as increasing the scale. Therefore, building the multiscale variant is not necessary with the sigmoid weighting.
@@ 4.2 Scaling channels vs tokens on ImageNet
In this section we show the difference between scaling a model 4x FLOPS, either via channels or by creating larger resolution feature maps (or put differently, feature maps with a higher token count). Note that our small model on ImageNet512 is patched 4 x 4. This model is scaled either by increasing channels by two (named model size 2x), or by reducing patching to 2 x 2 (named flop heavy, because FLOPS over params is much higher). Indeed as can be seen in Figure 10, the flop heavy model performs much better than the model size 2x, when trained in a standard way on ImageNet512. Note that unlike most experiments, in these experiments it was necessary to use very different optimal guidance settings, more details in Appendix B.
We confirm this is an issue related to overfitting, by training the model size 2x with distribution augmentation [18] the model size 2x variant performs much better, with an FID of 1.6 instead of 1.9 (latter in Figure 10). Still, without any augmentation beyond the standard horizontal flips, the flop heavy model achieves a score of 1.54.
Table 2: Literature Comparison on ImageNet on FID. Note that even though certain NFEs are high (e.g. 4100 for CDM) typically these suffer minor performance degradation for lower NFEs and should not be taken too literally.
Method NFE FID
@@
RIN [16] 1000 2.75
SD, U-ViT-L [14] 512 1.94
StyleGAN-XL [38] 1 1.81
VDM$+$ [25] 250 1.75
DisCo-Diff [46] 414 1.73
PaGoDA [22] 1 1.48
SiD2, Flop Heavy (ours) 512 1.26
@@
ADM-G [4] 250 4.59
ADM-U [4] 250 3.94
CDM [12] 4100 4.88
RIN [16, 3] 1000 3.52
SD, U-ViT-L [14] 512 2.44
MDM [7] 250 3.51
StyleGAN-XL [38] 1 2.30
DiT-XL/2 [31] 250 2.27
VDM++ [25] 250 2.12
RDM [42] 356 1.89
PaGoDA [22] 1 1.56
SiD2, small (ours) 512 1.72
SiD2, Flop Heavy (ours) 512 1.38
@@
ADM-G [4] 250 7.72
RIN [16, 3] 1000 3.95
ADM-U [4] 250 3.85
LDM-4-G [32] 250 3.60
SD, U-ViT-L [14] 512 3.02
DiffiT [9] 250 2.67
StyleGAN-XL [38] 1 2.40
DiT-XL/2 [31] 250 3.04
VDM++ [25] 250 2.65
MAGVIT-v2 [51] 64 1.91
PaGoDA [22] 1 1.80
EDM2-S (interval) [27] 63 1.68
EDM2-XXL (interval) [27] 63 1.40
SiD2, small (ours) 512 2.19
SiD2, Flop Heavy (ours) 512 1.48
SiD2, Flop Heavy distilled (ours) 16 1.50
Imagenet 1024 x 1024
RIN [16] 1000 8.72
SiD2, Flop Heavy (ours) 512 1.75
@@ 4.3 Removing block skip connection
In this section we study the difference between the residual skip-connections and the blockwise skip-connections. As hypothesized, the small model performance drops a bit because of this. For ImageNet512 a deficit of about 0.2 FID (in Figure  10) and for Imagenet256 a deficit of about 0.05 FID. That the task of fitting ImageNet256 is less affected by removing skip-connections can be explained by the task being easier: The task of fitting ImageNet256 is a subset of fitting ImageNet512.
@@
[…]
DeepFloyd [40] 256 6.7
eDiff-I [1] 256 6.9
Parti [49] 256 7.2
Imagen [35] 256 7.3
simple diffusion [14] 256 8.3
simple diffusion [14] 512 256 9.6
SDv1.5 (eval DMD2) [32] 512 200 7.2
DMD2 [48] 512 4 8.3
SiD2 (ours, bias=-3) 512 256 8.1
SiD2 16-step distilled (ours) 512 16 6.7
@@ 4.4 Literature comparison
Here we compare different models from the literature. In the first section end-to-end diffusion models are compared, and in the second section we compare between all types of generative models. In Table 2 it can be seen that our SiD2 is very competitive with existing generative models: On ImageNet128 and ImageNet256 it achieves state-of-the-art performance, on ImageNet512 it is a little worse but still competitive with the best latent diffusion model and outperforms all other models, pixel-based and latent.
@@
The training cost SiD2 is another helpful way to see how the model scales, as depicted in Figure 7). Importantly, the performance of our pixel-based approach outperforms other significantly. Observe for example the performance of our small variant (using no guidance interval to keep this comparison fair), which roughly the same performance as the VDM++ while only using a third of the training cost, and a fourth of the FLOPS in a single forward pass (see Figure 16 in Appendix C). It is also competitive with the best currently known latent diffusion approach EDM2 [20], but in fairness EDM2 scales a little bit better. These models attain even better performance for a smaller training cost, and a similar story holds for forward pass complexity. End-to-end diffusion can work much better than previously demonstrated, but it is still somewhat outperformed by latent diffusion on quality versus compute comparisons for both sampling and training.
Table 4: Conditional Video generation performance on Kinetics600 in FVD (resized to 64). For video generation the loss specification is important to obtain good sample quality (see ablations).
[…]
pixel diffusion
Video Diffusion Models [13] 16.6
RIN [16] 10.8
Rolling Diffusion [34] 5.2
latent diffusion
MAGVIT [50] 9.9
MAGVITv2 [51] 4.3
W.A.L.T [8] 3.3
SiD2 (ours, sigmoid bias = 0) 2.3
ablations
        eps-mse 5.7
        sigmoid bias = +2 3.2
        sigmoid bias = -2 2.5
@@
Video Generation
To test whether our approach transfers to video, we test our model on Kinetics600 following [13], by conditioning on 5 frames and generating (sometimes referred to as predicting) the subsequent 11 frames. We then compute the FVD between the training set and the generated samples both on the subsequent 11 frames. Even though our method generates at $128\times 128$, we follow [8] and resize samples to have resolution $64\times 64$. Our model achieves SOTA performance and it turns out that tuning the loss is very important to achieve this as shown in Table 4. Note that the latent approaches ([50, 51, 8]) use much less training compute than our models in pixel-space, which we discuss in more detail in the Appendix B.